# The Active Optics System on the Vera C. Rubin Observatory: Optimal Control of Degeneracy Among the Large Number of Degrees of Freedom

### v1

@@ Abstract
The Vera C. Rubin Observatory is a unique facility for survey astronomy that will soon be commissioned and begin operations. Crucial to many of its scientific goals is the achievement of sustained high image quality, limited only by the seeing at the site. This will be maintained through an Active Optics System (AOS) that controls optical element misalignments and corrects mirror figure error to minimize aberrations caused by both thermal and gravitational distortions. However, the large number of adjustment degrees of freedom available on the Rubin Observatory introduces a range of degeneracies, including many that are noise-induced due to imperfect measurement of the wavefront errors. We present a structured methodology for identifying these degeneracies through an analysis of image noise level. We also present a novel scaling strategy based on Truncated Singular Value Decomposition (TSVD) that mitigates the degeneracy, and optimally distributes the adjustment over the available degrees of freedom. Our approach ensures the attainment of optimal image quality, while avoiding excursions around the noise-induced subspace of degeneracies, marking a significant improvement over the previous techniques adopted for Rubin, which were based on an Optimal Integral Controller (OIC). This new approach is likely to also yield significant benefits for all telescopes that incorporate large numbers of degrees of freedom of adjustment.
@@ 1 Introduction
The forefront ground-based wide-field survey facility, the Vera C. Rubin Observatory, will soon begin on-sky commissioning on Cerro Pachón in Chile. It is poised to address a variety of pressing scientific questions over its 10-year Legacy Survey of Space and Time (LSST) (Ivezić et al., 2019). The Rubin Observatory incorporates the Simonyi Survey Telescope, an 8.4m-diameter modified Paul-Baker three-mirror anastigmat, with a remarkable $3.5^{\circ}$ field of view (FOV). To deliver the required seeing-limited image resolution across its field, the tolerance on image distortions introduced by the telescope and camera is limited to 0.4 arcsec (Claver et al., 2023). This is planned to be achieved through an Active Optics System (AOS) that uses wavefront sensing to control a large-dimensional parameter space of alignment and figure adjustments on each of the three mirrors, as well as the translation and rotation of the camera (Neill et al., 2014).
@@ 3 Active Optics
The AOS on Rubin is designed to correct the wavefront distortions induced by gravitational and thermal effects in real-time. This involves sensing aberrations in the wavefront from images and subsequently adjusting the telescope’s degrees of freedom, including surface figure modes and misalignments of the optical elements. Typically, an AOS comprises both open and closed loop components, collectively working to minimize the wavefront deviation.
@@ 3.1 Misalignments and Mirror Bending Modes
Table 3: Available mirror bending mode ranges derived by uniformly distributing the actuator force range across all bending modes to ensure equal force allocation per mode. The range per mode is then calculated from this allocated force using the influence matrix that relates actuator forces (N) to bending modes ($\mu$m).
@@ 3.3 Closed loop component
Once the estimate of the wavefront deviation is obtained, the sensitivity matrix is “inverted” to determine the estimated corrections to the degrees of freedom. In the prior Rubin AOS software system, the final applied corrections were subsequently derived through an OIC controller (Neill et al., 2014; Homar, 2024). The formulation and details of this process are discussed in the following section, along with the exploration of degeneracies in both state estimation and correction calculations.
@@ 6.1 Characteristic modes
Lastly, mode $\tilde{v}_{40}$ in the rescaled basis, combining $B_{1,6}$, $B_{1,15}$, and to a lesser extent $B_{1,13}$, $B_{1,17}$, and $B_{2,7}$, also warrants mention. The bending mode Figures 1 and 2 suggest why these combinations minimally affect or neutralize each other’s impact on the wavefront. A similar rationale, grounded in optical intuition, can be applied with varying degrees of ease to the remainder of the characteristic modes to discern the interplay among different degrees of freedom.
Examining these results, it becomes clear the characteristic modes do not necessarily represent the actual state of the telescope’s optics but instead indicate the ”easiest” way, within the system’s limitations, of producing the measured wavefront errors. This will not guarantee the precision of the estimated optical state; however, the contributions from the estimated state will provide a comparable wavefront and are the best estimate and strategy for correction we can derive. For example, in the presence of noise, it’s impossible to determine whether the optical aberration caused by $B_{1,1}$ originates from M1M3 or its counterpart $B_{2,2}$ on M2. This ambiguity underscores the value of characteristic modes as a crucial framework for estimating the optical state and implementing system corrections. Indeed, the rescaled characteristic mode basis selects M2 as the prevalent mode, reflecting its greater corrective capacity within the system.
Finally, it is worth mentioning the changes observed in the singular value distribution. For the rescaled matrix we still have significantly lower singular values, but the decay is steeper, an effect arising from the rescaling of the different modes.
@@ 6.2 Noise effects
The manifestation of noise becomes pronounced from mode $\tilde{u}_{40}$ onwards for two primary reasons. Modes beyond $\tilde{u}_{50}$ are exclusively included because of SVD basis completion which extends the basis from 50 $u_{i}$ singular vectors up to 76 through the null space of the sensitivity matrix, they are thus absent in our measurements and bear null singular values. Modes between $\tilde{u}_{40}$ and $\tilde{u}_{50}$ correspond to those with the lowest singular values in Figure 5, with minimal impact on the wavefront. However, the onset of noise appears to be close to $\tilde{u}_{27}$, with the first noise peak being observed at $\tilde{u}_{30}$, which corresponds in Figure 5 to a characteristic mode involving a quasi-degenerate counter-movement of pistoning between M2 and the Camera.
This analysis, in conjunction with the right panel’s cumulative singular value sum, guides our selection of the truncation threshold. The power spectrum suggests mode 27 as a plausible truncation point, a conclusion bolstered by the cumulative sum indicating that 99% of the singular value contribution is reached by the 21st mode ($\tilde{\sigma}_{21}$). Despite this, to maximize the information gleaned from the wavefront measurement, we extend the truncation to the 27th mode.
@@ 6.3 Controller
The OIC method achieves image quality convergence but fails to suppress movements within the noise-induced degenerate subspace. This limitation arises because the current OIC control strategy lacks a regularization term that effectively penalizes the movements in this subspace. While future work will explore different penalty possibilities, these will still to be founded on the degenerate subspace that we have identified here.
@@ 7 Discussion and Conclusion
We would like to thank Roberto Tighe and Aaron Roodman for helpful discussions on mirror prioritization, Tiago Ribeiro for discussions on the current baseline, Doug Niell and Ellie Hilerman for clarifications on the degree of freedom ranges and mirror flexures, and Andy Rasmussen and Kevin Reil for reviewing this manuscript. This material is based upon work supported in part by the National Science Foundation through Cooperative Agreement AST-1258333 and Cooperative Support Agreement AST-1202910 managed by the Association of Universities for Research in Astronomy (AURA), and the Department of Energy under Contract No. DE-AC02-76SF00515 with the SLAC National Accelerator Laboratory managed by Stanford University. Additional Rubin Observatory funding comes from private donations, grants to universities, and in-kind support from LSST-DA Institutional Members. The project leading to these results also received funding from “la Caixa” Foundation (ID 100010434), under the agreement LCF/BQ/EU21/11890114. This research made use of Stanford University’s computational resources. G.M. thanks the LSSTC Data Science Fellowship Program, which is funded by LSSTC, NSF Cybertraining Grant #1829740, the Brinson Foundation, and the Moore Foundation; his participation in the program has benefited this work. J.F.C. acknowledges support from the U.S. Department of Energy, Office of Science, Office of High Energy Physics Cosmic Frontier Research program under Award Number DE-SC0011665. The Vera C. Rubin 8.4m Simonyi Survey Telescope Observatory

### v2

@@ Abstract
The Vera C. Rubin Observatory is a unique facility for survey astronomy that will soon be commissioned and begin operations. Crucial to many of its scientific goals is the achievement of sustained high image quality, limited only by the seeing at the site. This will be maintained through an Active Optics System (AOS) that controls optical element misalignments and corrects mirror figure error to minimize aberrations caused by both thermal and gravitational effects. However, the large number of adjustment degrees of freedom available on the Rubin Observatory introduces a range of degeneracies, including many that are noise-induced due to imperfect measurement of the wavefront errors. We present a structured methodology for identifying these degeneracies through an analysis of image noise level. We also present a novel scaling strategy based on Truncated Singular Value Decomposition (TSVD) that mitigates the degeneracy, and optimally distributes the adjustment over the available degrees of freedom. Our approach ensures the attainment of optimal image quality, while avoiding excursions around the noise-induced subspace of degeneracies, marking a significant improvement over the previous techniques adopted for Rubin, which were based on an Optimal Integral Controller (OIC). This new approach is likely to also yield significant benefits for all telescopes that incorporate large numbers of degrees of freedom of adjustment.
@@ 1 Introduction
The forefront ground-based wide-field survey facility, the Vera C. Rubin Observatory, will soon begin on-sky commissioning on Cerro Pachón in Chile. It is poised to address a variety of pressing scientific questions over its 10-year Legacy Survey of Space and Time (LSST) (Ivezić et al., 2019). The Rubin Observatory incorporates the Simonyi Survey Telescope, an 8.4m-diameter modified Paul-Baker three-mirror anastigmat, with a remarkable $3.5^{\circ}$ field of view (FOV). To deliver the required seeing-limited image resolution across its field, the tolerance on image aberrations introduced by the telescope and camera is limited to 0.4 arcsec (Claver et al., 2023). This is planned to be achieved through an Active Optics System (AOS) that uses wavefront sensing to control a large-dimensional parameter space of alignment and figure adjustments on each of the three mirrors, as well as the translation and rotation of the camera (Neill et al., 2014).
@@ 3 Active Optics
The AOS on Rubin is designed to correct the wavefront aberrations induced by gravitational and thermal effects in real-time. This involves sensing aberrations in the wavefront from images and subsequently adjusting the telescope’s degrees of freedom, including surface figure modes and misalignments of the optical elements. Typically, an AOS comprises both open and closed loop components, collectively working to minimize the wavefront deviation.
@@ 3.1 Misalignments and Mirror Bending Modes
Table 3: Available mirror bending mode ranges derived by uniformly distributing the actuator force range across all bending modes to ensure equal force allocation per mode. The range per mode is calculated from this allocated force using the influence matrix that relates actuator forces (N) to bending modes ($\mu$m in surface deviation).
@@ 3.3 Closed loop component
Once the estimate of the wavefront deviation is obtained, the sensitivity matrix is “inverted” to determine the estimated corrections to the degrees of freedom. In the prior Rubin AOS software system, the final applied corrections were subsequently derived through an OIC controller (Neill et al., 2014; Megias Homar, 2024). The formulation and details of this process are discussed in the following section, along with the exploration of degeneracies in both state estimation and correction calculations.
@@ 6.1 Characteristic modes
Lastly, mode $\tilde{v}_{43}$ in the rescaled basis, combining $B_{1,6}$, $B_{1,15}$, and to a lesser extent $B_{1,13}$, $B_{1,17}$, and $B_{2,7}$, also warrants mention. The bending mode Figures 1 and 2 suggest why these combinations minimally affect or neutralize each other’s impact on the wavefront. A similar rationale, grounded in optical intuition, can be applied with varying degrees of ease to the remainder of the characteristic modes to discern the interplay among different degrees of freedom.
Examining these results, it becomes clear the characteristic modes do not necessarily represent the actual state of the telescope’s optics but instead indicate the “easiest” way, within the system’s limitations, of producing the measured wavefront errors. This will not guarantee the precision of the estimated optical state; however, the contributions from the estimated state will provide a comparable wavefront and are the best estimate and strategy for correction we can derive. For example, in the presence of noise, it’s impossible to determine whether the optical aberration caused by $B_{1,1}$ originates from M1M3 or its counterpart $B_{2,2}$ on M2. This ambiguity underscores the value of characteristic modes as a crucial framework for estimating the optical state and implementing system corrections. Indeed, the rescaled characteristic mode basis selects M2 as the prevalent mode, reflecting its greater corrective capacity within the system.
Finally, it is worth mentioning the changes observed in the singular value distribution. For the rescaled matrix we still have significantly lower singular values, but the decay is steeper, effectively resembling a power law, an effect arising from the rescaling of the different modes.
@@ 6.2 Noise effects
The manifestation of noise becomes pronounced from mode $\tilde{u}_{40}$ onwards for two primary reasons. Modes beyond $\tilde{u}_{50}$ are exclusively included because of SVD basis completion which extends the basis from 50 $u_{i}$ singular vectors up to 76 through the null space of the sensitivity matrix, they are thus absent in our measurements and bear null singular values. Modes between $\tilde{u}_{40}$ and $\tilde{u}_{50}$ correspond to those with the lowest singular values in Figure 5, with minimal impact on the wavefront. However, the onset of noise appears to be close to $\tilde{u}_{25}$, with the first noise peak being observed at $\tilde{u}_{30}$, which corresponds in Figure 5 to a characteristic mode involving a quasi-degenerate counter-movement of pistoning between M2 and the Camera.
This analysis, in conjunction with the right panel’s cumulative singular value sum, guides our selection of the truncation threshold. The power spectrum suggests mode 25 as a plausible truncation point, a conclusion bolstered by the cumulative sum indicating that 99% of the singular value contribution is reached by the 25th mode ($\tilde{\sigma}_{25}$).
@@ 6.3 Controller
The OIC method achieves image quality convergence but fails to suppress movements within the noise-induced degenerate subspace. This limitation arises because the current OIC control strategy lacks a regularization term that effectively penalizes the movements in this subspace, since it penalizes the incremental motion instead of the total optical state. While future work will explore different penalty possibilities, these will still to be founded on the degenerate subspace that we have identified here.
@@ 7 Discussion and Conclusion
The rescaling of the sensitivity matrix ensures the suppression of degenerate motions by embedding system-specific design constraints and priorities, resulting in optimized performance. Interestingly, our approach is adaptable and can be applied to various systems and different design constraints, yielding a different non-degenerate basis for each configuration.
@@
We would like to thank Roberto Tighe and Aaron Roodman for helpful discussions on mirror prioritization, Tiago Ribeiro for discussions on the current baseline, Doug Niell and Ellie Hilerman for clarifications on the degree of freedom ranges and mirror flexures, Andy Rasmussen and Kevin Reil for reviewing this manuscript, and Will Sutherland for his optical intuition insights on an earlier version of this paper. This material is based upon work supported in part by the National Science Foundation through Cooperative Agreement AST-1258333 and Cooperative Support Agreement AST-1202910 managed by the Association of Universities for Research in Astronomy (AURA), and the Department of Energy under Contract No. DE-AC02-76SF00515 with the SLAC National Accelerator Laboratory managed by Stanford University. Additional Rubin Observatory funding comes from private donations, grants to universities, and in-kind support from LSST-DA Institutional Members. The project leading to these results also received funding from “la Caixa” Foundation (ID 100010434), under the agreement LCF/BQ/EU21/11890114. This research made use of Stanford University’s computational resources. G.M. thanks the LSSTC Data Science Fellowship Program, which is funded by LSSTC, NSF Cybertraining Grant #1829740, the Brinson Foundation, and the Moore Foundation; his participation in the program has benefited this work. J.F.C. acknowledges support from the U.S. Department of Energy, Office of Science, Office of High Energy Physics Cosmic Frontier Research program under Award Number DE-SC0011665. The Vera C. Rubin 8.4m Simonyi Survey Telescope Observatory